9.3 a) For the beam shown in Fig. 9.3, draw the shear force and bending moment diagrams. Use the area method that relies on tI just need A in detail
Answer #1

Calculation for sf 6RN 4 RN fra MP tam-2. 0D dm ac KRP Efuzo Rat RA = 6 + 4x2 - 14 KN . EMA=0 -4x285 – 6x2 + MAZO MA= 52 kNm.& Sof jumps o sf down by 6 kN. just to the night of B = 14-6= 8 kort e B to c There is no loading de =0 ie vis constant C toAt A BM = -52. KN-ne cine means hogging. co A to B dH = V(shear а и Ino force) slope of Bi Bч) 7 diagrame and constant and coan OM = V = slope M BMD, Heuse slope is positive but decreasing AM = Mo - Me = ama o SAD blwoande ro - (-89 = 2 & 8x2 t que B